Operating Activities Fiscal year is September-August. All values USD millions. 2014 2015 2016 2017 2018 Net Income before Extraordinaries 2.09B 2.41B 2.38B 2.71B 3.18B Depreciation, Depletion & Amortization 1.03B 1.13B 1.26B 1.37B 1.44B Depreciation and Depletion Amortization of Intangible Assets Deferred Taxes & Investment Tax Credit (63M) (101M) 269M (29M) (49M) Deferred Taxes (63M) (101M) 269M (29M) (49M) Investment Tax Credit 402M Other Funds Funds from Operations Extraordinaries Changes in Working Capital 265M 303M 500M 538M 3.32B 3.74B 4.3B 4.56B 5.11B 665M 547M (1.01B) 2.17B 669M Receivables 529M 699M 3.988 (1.53B) 547M 3.29B 1.56B 421M 5.77B Accounts Payable 880M 557M 4.29B 2.26B 807M 6.738 Other Assets/Liabilities Net Operating Cash Flow

Using the above operating activities for Costco Corporation, answer the following question:

Compare the net cash provided by operations with the net income for each of the years presented in the statement of cash flows above. How do the net cash flows from operating activities compare to the net income? Have the net cash flows from operating activities been increased or decreased over time?

Thanks!
Answer #1

There's a useful ratio we can deploy against the income statement to determine the quality of a firm's reported earnings: Operating Cash Flow (CFFO) to Net Income.

As the name implies, the ratio is calculated by dividing a company's operating cash flow by its net income. Operating activities are the actions a company takes pursuant to its normal business activities.

When a company's CFFO to net income ratio is above 1, it is indicative of a strong ability to fund it's activities through generation of operating cash flow. In other words, a higher ratio means that the firm's earnings are of a higher quality, and that it will not need to raise money any time soon to fund its operations.

Here are the calculations for the CFFO to Net Income Ratio:

2014: 3.98/2.09 = 1.90

2015: 4.29/2.41 = 1.78

2016: 3.29/2.38 = 1.38

2017: 6.73/2.71 = 2.48

2018: 5.77/3.18 = 1.81

Thus, the ratio has fluctuated downwards for the first three years. 2017 seems to have been a very good year for the company in terms of operations and the earnings quality was very high. That ratio seems to have normalized again in 2018.

The net cash flow from operating activities have fluctuated greatly in the period shown, increasing from 2014 to 2015, then reducing in 2016, before bouncing back up again strongly in 2017 and then normalizing in 2018 to a healthy level.